# When Truth Is Overridden: Uncovering the Internal Origins of Sycophancy in Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) often exhibit sycophantic behavior, agreeing with user-stated opinions even when those contradict factual knowledge. While prior work has documented this tendency, the internal mechanisms that enable such behavior remain poorly understood. In this paper, we provide a mechanistic account of how sycophancy arises within LLMs. We first systematically study how user opinions induce sycophancy across different model families. We find that simple opinion statements reliably induce sycophancy, whereas user expertise framing has a negligible impact. Through logit-lens analysis and causal activation patching, we identify a two-stage emergence of sycophancy: (1) a late-layer output preference shift and (2) deeper representational divergence. We also verify that user authority fails to influence behavior because models do not encode it internally. In addition, we examine how grammatical perspective affects sycophantic behavior, finding that first-person prompts (“I believe…”) consistently induce higher sycophancy rates than third-person framings (“They believe…”) by creating stronger representational perturbations in deeper layers. These findings highlight that sycophancy is not a surface-level artifact but emerges from a structural override of learned knowledge in deeper layers, with implications for alignment and truthful AI systems.
@@ Introduction
Figure 1: Overview of prompt types in experiments. More examples and details can be found in Appendix.
@@
Alignment techniques for Large Language Models (LLMs), such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022) and Direct Preference Optimization (DPO) (Rafailov et al. 2023), are widely employed to better align model behavior with human expectations and values (Wang et al. 2023). However, recent studies have revealed a critical drawback: LLMs with or without certain alignment techniques can inadvertently promote “sycophancy” (Casper et al. 2023), a behavior where models generate responses that cater to user beliefs or expectations, even when these deviate from truth (Sharma et al. 2023; Denison et al. 2024). This issue gained public attention especially after the April 2025 rollback of OpenAI’s GPT-4o (OpenAI 2025), which was widely condemned for uncritically mirroring user sentiments, regardless of their accuracy or potential for harm.
@@
Prior work has extensively documented this sycophantic behavior across model sizes and training paradigms (Perez et al. 2023), developing intervention methods using synthetic data, steering vectors, pinpoint tuning, and DPO to successfully reduce such responses (Wei et al. 2023; Panickssery et al. 2023; Chen et al. 2024; Khan et al. 2024). However, these approaches primarily focus on controlling the behavior rather than understanding the underlying mechanism. This gap between behavioral control and mechanistic understanding motivates our investigation into how sycophancy manifests within model computations. Recent studies have shown that language models can be influenced by user inputs that contain opinions or statements that contradict the model’s learned knowledge (Sharma et al. 2023; Fanous et al. 2025). Hence, we need to examine the computational mechanisms that occur when models process such conflicting information, as illustrated in Figure 1. In this paper, we trace how sycophantic behavior emerges through the model’s architecture and analyze the stages of processing where user opinions begin to override learned knowledge. Our central questions focus on understanding where and how this representational shift occurs, and what specific mechanisms allow user opinion framing to influence the model’s final outputs, even when contradicting information that the model has previously learned.
@@
To achieve this goal, we first designed a straightforward experimental framework that avoids the complexity of existing multi-stage benchmarks and subjective LLM-as-a-judge evaluations. Following prior research showing that incorrect user opinions can reliably trigger sycophantic behavior in many cases (Sharma et al. 2023; Fanous et al. 2025), we use this as our primary sycophancy trigger across seven model families of similar size, with MMLU (Hendrycks et al. 2020) as our dataset for its multi-subject coverage and multiple-choice format. Moreover, we extended the basic opinion-based approach by incorporating three levels of perceived user expertise: Beginner, Intermediate, Advanced. This design allows us to distinguish between two potential mechanisms: opinion-driven sycophancy (models conform simply because users express opinions) versus authority-driven sycophancy (models are additionally influenced by perceived user credibility). Our results show that simple user opinions (“I believe the right answer is…”) consistently induce sycophancy across all seven models, while different expertise levels do not significantly affect sycophancy rates.
@@
Based on our results, we then analyzed the phenomenon of internal sycophantic space from a mechanistic perspective. We find that user opinions prevent the emergence of fact-based preferences that would otherwise develop in later layers, as evidenced by logit-lens (nostalgebraist 2020) analysis and validated through causal activation patching (Wang et al. 2022) experiments, where interventions at critical layers can reduce sycophancy. As for why expertise levels do not significantly modulate sycophancy, we examined how models internally represent users with different expertise levels. We found that the representations largely overlap rather than form distinct patterns, indicating that models fail to encode user expertise as a meaningful factor in their processing.
@@
We also tested whether the way users express their own opinions matters by comparing direct statements (“I believe…”) with indirect ones (“They believe…”) as illustrated in Figure 1. Inspired by research on how models respond to indirect, third-person suggestions (Chen et al. 2025) and cognitive science findings showing that people are less influenced by third-person versus first-person perspectives in social conformity (Wallace-Hadrill and Kamboj 2016), we examined this perspective-driven effect across all models. We found that indirect third-person statements consistently reduce sycophancy compared to direct first-person statements.
@@
Prior work established that sycophancy scales with model size and appears across training paradigms (Perez et al. 2023). Research on RLHF revealed a mechanism: models can prioritize user satisfaction over factual accuracy through reward hacking, learning to maximize human approval rather than truthfulness (Stiennon et al. 2020). Sharma et al. (2023) investigated why this happens by analyzing the training data itself, discovering that human preference datasets contain inherent biases that teach models to agree with users rather than provide accurate information, leading to benchmarks such as SycEval (Fanous et al. 2025), a multi-round evaluation framework to measure and categorize sycophancy.
Recent work has expanded our understanding of sycophancy from different perspectives. Cheng et al. (2025) identified a form called “social sycophancy”, where models avoid providing feedback that might hurt users’ feelings or self-image. Meanwhile, (Zhao et al. 2024) demonstrated that sycophantic patterns emerge in vision-language models when processing visual content alongside user commentary.
Efforts to reduce sycophancy have provided insights into its underlying mechanisms. Synthetic data interventions can teach models to resist user pressure and maintain factual accuracy (Wei et al. 2023), while targeted fine-tuning like pinpoint tuning addresses internal representations that drive sycophantic responses (Chen et al. 2024). Sicilia, Inan, and Alikhani (2024) addressed how models inappropriately mirror user confidence levels, developing uncertainty-based methods to help models express their own epistemic doubt. Other work explored steering vectors and contrastive activation methods to manipulate activations responsible for sycophancy (Panickssery et al. 2023), revealing that sycophancy emerges from specific neural activity patterns that can be identified and modified.
@@
Mechanistic interpretability (MI) aims to reverse-engineer neural networks into human-interpretable algorithms, moving beyond traditional explainable AI focused on input-output relationships (Bereska and Gavves 2024; Kästner and Crook 2024). Key techniques for transformers include logit-lens analysis (nostalgebraist 2020), which extracts meaningful token predictions from intermediate layers, revealing what the model “believes” after each step and how these distributions converge toward the final output (Stolfo, Belinkov, and Sachan 2023). Activation patching provides a causal perspective by substituting activations between inputs to identify which components are necessary and sufficient for specific behaviors (Wang et al. 2022).
Recent applications of these MI techniques to sycophantic behavior have begun to illuminate the internal mechanisms involved, though with limitations. Yu, Merullo, and Pavlick (2023) used head attribution to identify attention heads that resolve conflicts between memorized facts and contradictory contextual information, but their findings on factual recall (e.g., world capitals) show limited generalizability across different knowledge domains. Similarly, steering vector approaches have identified specific activation space directions corresponding to sycophantic versus truthful responding and can successfully modify model behavior through targeted interventions (Panickssery et al. 2023), but these methods focus on controlling sycophancy rather than explaining why these particular directions emerge or what computational processes give rise to the observed activation patterns.
@@
Following previous studies, “sycophancy” in LLMs is defined as the model’s tendency to conform to a user’s explicitly stated opinion, even when that opinion is incorrect (Denison et al. 2024). To understand how models process conflicting information when user opinions contradict learned knowledge, we designed a simple experimental framework that avoids multi-stage benchmark complexity (Fanous et al. 2025). Prior work shows that incorrect user opinions reliably trigger sycophancy (Sharma et al. 2023), so we use this as our primary sycophancy trigger across seven model families.
@@
Models. We evaluate seven LLMs from different families: Llama3.1 8B-Instruct (Dubey et al. 2024), Qwen2.5 7B-Instruct (Team 2024a, b), OPT 6.7B (Zhang et al. 2022), Mistral 7B-Instruct v0.3 (Jiang et al. 2023), Falcon 7B (Technology Innovation Institute 2023), OLMoE 1B-7B Instruct (Muennighoff et al. 2024), and Pythia 6.9B (Biderman et al. 2023), all of comparable parameter size, ensuring observed differences are not due to scale effects.
Dataset. We use the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al. 2020). Its multiple-choice format provides clear ground-truth labels, while its 57 academic subjects ensure broad generalizability across knowledge domains. In our case, sycophancy occurs when the model selects the same answer choice that the user claims to believe, which we designed to always be incorrect.
@@
Evaluation Metric. For each sample, we log the model’s selected answer, and then compute three metrics: (1) sycophancy rate, or agreement rate (Malmqvist 2024), the proportion of samples where the model agrees with the user’s belief; (2) accuracy, the proportion where it selects the correct answer; and (3) independent error rate, the proportion of incorrect answers that disagree with both the user’s belief and the ground truth, indicating autonomous errors.
@@
Figure 3: Responses breakdown by user expertise level. Results show expertise has negligible impact on sycophancy rates. A detailed table is at Table 2 in Appendix.
@@
Our first objective is to identify when during the model’s forward pass sycophantic preferences emerge. Since different transformer layers encode different types of information, with later layers typically handling task-specific reasoning (Tenney, Das, and Pavlick 2019; Rogers, Kovaleva, and Rumshisky 2021), we hypothesize that sycophancy arises at a specific computational stage where user opinion overrides LLM’s learned knowledge.
To test this, we design Decision Score, a layer-wise metric designed to track how the model’s internal preference shifts between the correct answer and the user’s stated (incorrect) opinion. At each layer of the transformer, we take the model’s hidden states (internal representations) to predict which answer it would choose if it stopped there. This lets us see how its preference changes as information flows through the network. Doing this gives us the prediction scores (called logits) for each of the four multiple-choice options: $l_{A},l_{B},l_{C},l_{D}$ via logit-lens (nostalgebraist 2020) (more details in Appendix). For any candidate option $x\in\{A,B,C,D\}$, we define a normalized score:
@@
Figure 4: Layered decision scores of the correct and chosen wrong answers under Plain and Opinion-only on Llama3.1 8B-Instruct. Result for Qwen can be found in Figure 11 in Appendix.
@@
Having identified when sycophancy emerges, we next investigate how opinion framing alters the model’s internal representations. Prior work has demonstrated that different prompting strategies can lead to measurably different activation patterns (Wei et al. 2022; Hendel, Geva, and Globerson 2023), suggesting that semantic framing effects should be detectable in hidden state distributions.
@@
While the above methods reveal correlations, establishing causality requires direct intervention. Activation patching tests whether specific internal changes are necessary and sufficient for a behavior (Meng et al. 2022; Yeo, Satapathy, and Cambria 2024; Zhang and Nanda 2023). If our observed representational shifts truly drive sycophancy, then selectively modifying these activations should predictably alter the output.
Figure 6: Causal effects of activation patching, for both models, at respective turning-point layers found in KL divergence. (a) Plain $\rightarrow$ Opinion-only : Patching induces sycophancy (orange ↑) while reducing accuracy (green ↓); (b) Opinion-only $\rightarrow$ Plain : Patching suppresses sycophancy (orange ↓) and improves accuracy (green ↑).
@@
Figure 7: PCA projection of prompt token hidden states from Layer 32 of Llama3.1 8B-Instruct, across four user framings: Opinion-only (light blue), First-pov with Beginner (red), Intermediate (green), and Advanced (blue). Result for Qwen cna be found at Figure 12 in Appendix.
@@
To investigate this, we turn to cognitive science, where research shows that narrative point-of-view fundamentally shapes human perception (Wallace-Hadrill and Kamboj 2016). A first-person perspective is associated with subjective, emotionally resonant experiences, whereas a third-person view fosters objectivity and psychological distance. Since LLMs learn from human-generated text, they may have implicitly learned to differentiate these frames. We therefore hypothesize that framing a belief in the third person will reduce sycophantic behavior compared to the First-pov prompts used in our initial experiments.
@@
Figure 8: 7 models consistently exhibit more sycophancy in First-pov than in Third-pov , with an average incrase of 13.6%, on all expertise levels (full result at Figure 14 in Appendix).
As shown in Figure 8, our experimental results reveal a consistent pattern: First-pov induces higher sycophancy rates than Third-pov . A plausible explanation of this behaviour, rooted in the cognitive science principles discussed previously (Wallace-Hadrill and Kamboj 2016), is that the first-person pronoun “I” is interpreted by the model as a direct, subjective appeal from the user. In contrast, the third-person “they” frames the belief as a detached, objective report about another entity. This psychological distance appears to reduce the pressure to conform, allowing the model’s internal knowledge to influence its final output more freely.
@@
To find mechanistic evidence for the behavioral differences observed, we investigated where the model’s internal representations diverge under these different person framings. Specifically, using layer-wise KL divergence, we measured the difference between the hidden state distributions of First-pov and Third-pov prompts against a Plain baseline.

### v3

@@ Abstract
Large Language Models (LLMs) often exhibit sycophantic behavior, agreeing with user-stated opinions even when those contradict factual knowledge. While prior work has documented this tendency, the internal mechanisms that enable such behavior remain poorly understood. In this paper, we provide a mechanistic account of how sycophancy arises within LLMs. We first systematically study how user opinions induce sycophancy across different model families. We find that simple opinion statements reliably induce sycophancy, whereas user expertise framing has a negligible impact. Through logit-lens analysis and causal activation patching, we identify a two-stage emergence of sycophancy: (1) a late-layer output preference shift and (2) deeper representational divergence. We also verify that user authority fails to influence behavior, because models do not encode it internally. In addition, we examine how grammatical perspective affects sycophantic behavior, finding that first-person prompts (“I believe…”) consistently induce higher sycophancy rates than third-person framings (“They believe…”) by creating stronger representational perturbations in deeper layers. These findings highlight that sycophancy is not a surface-level artifact but emerges from a structural override of learned knowledge in deeper layers, with implications for alignment and truthful AI systems.
Codegithub.com/kaustpradalab/LLM-sycophancy
@@ Introduction
Figure 1: Overview of prompt types in experiments. More examples and details can be found in the Appendix.
@@
Alignment techniques for Large Language Models (LLMs), such as Reinforcement Learning from Human Feedback (RLHF) (ouyang2022training) and Direct Preference Optimization (DPO) (rafailov2023direct), are widely employed to better align model behavior with human expectations and values (wang2023aligning; li2025curriculum; zhang2025towards; jin2025towards; kong2024perplexity). However, recent studies have revealed a critical drawback: LLMs with or without certain alignment techniques can inadvertently promote “sycophancy” (casper2023open), a behavior where models generate responses that cater to user beliefs or expectations, even when these deviate from truth (sharma2023towards; denison2024sycophancy; guo2025benchmarking; zhou2025flattery). This issue gained public attention especially after the April 2025 rollback of OpenAI’s GPT-4o (openai_sycophancy_4o_2025), which was widely condemned for uncritically mirroring user sentiments, regardless of their accuracy or potential for harm.
@@
Prior work has extensively documented this sycophantic behavior across model sizes and training paradigms (perez2023discovering), developing intervention methods using synthetic data, steering vectors, pinpoint tuning, and DPO to successfully reduce such responses (wei2023simple; panickssery2023steering; chen2024yes; khan2024mitigating). However, these approaches primarily focus on controlling the behavior rather than understanding the underlying mechanism. This gap between behavioral control and mechanistic understanding motivates our investigation into how sycophancy manifests within model computations. Recent studies have shown that language models can be influenced by user inputs that contain opinions or statements that contradict the model’s learned knowledge (sharma2023towards; fanous2025syceval; yang2024makes). Hence, we need to examine the computational mechanisms that occur when models process such conflicting information, as illustrated in Figure 1. In this paper, we trace how sycophantic behavior emerges through the model’s architecture and analyze the stages of processing where user opinions begin to override learned knowledge. Our central questions focus on understanding where and how this representational shift occurs, and what specific mechanisms allow user opinion framing to influence the model’s final outputs, even when contradicting information that the model has previously learned.
@@
To achieve this goal, we first designed a straightforward experimental framework that avoids the complexity of existing multi-stage benchmarks and subjective LLM-as-a-judge evaluations. Following prior research showing that incorrect user opinions can reliably trigger sycophantic behavior in many cases (sharma2023towards; fanous2025syceval), we use this as our primary sycophancy trigger across seven model families of similar size, with MMLU (hendrycks2020measuring) as our dataset for its multi-subject coverage and multiple-choice format. Moreover, we extended the basic opinion-based approach by incorporating three levels of perceived user expertise: Beginner, Intermediate, Advanced. This design allows us to distinguish between two potential mechanisms: opinion-driven sycophancy (models conform simply because users express opinions) versus authority-driven sycophancy (models are additionally influenced by perceived user credibility). Our results show that simple user opinions (“I believe the right answer is…”) consistently induce sycophancy across all seven models, while different expertise levels do not significantly affect sycophancy rates.
@@
Based on our results, we then analyzed the phenomenon of internal sycophantic space from a mechanistic perspective. We find that user opinions prevent the emergence of fact-based preferences that would otherwise develop in later layers, as evidenced by logit-lens (nostalgebraist2020interpreting) analysis and validated through causal activation patching (wang2022interpretability) experiments, where interventions at critical layers can reduce sycophancy. As for why expertise levels do not significantly modulate sycophancy, we examined how models internally represent users with different expertise levels. We found that the representations largely overlap rather than form distinct patterns, indicating that models fail to encode user expertise as a meaningful factor in their processing.
@@
We also tested whether the way users express their own opinions matters by comparing direct statements (“I believe…”) with indirect ones (“They believe…”) as illustrated in Figure 1. Inspired by research on how models respond to indirect, third-person suggestions (chen2025reasoning) and cognitive science findings showing that people are less influenced by third-person versus first-person perspectives in social conformity (wallace2016impact), we examined this perspective-driven effect across all models. We found that indirect third-person statements consistently reduce sycophancy compared to direct first-person statements.
@@
Prior work established that sycophancy scales with model size and appears across training paradigms (perez2023discovering). Research on RLHF revealed a mechanism: models can prioritize user satisfaction over factual accuracy through reward hacking, learning to maximize human approval rather than truthfulness (stiennon2020learning). sharma2023towards investigated why this happens by analyzing the training data itself, discovering that human preference datasets contain inherent biases that teach models to agree with users rather than provide accurate information, leading to benchmarks such as SycEval (fanous2025syceval), a multi-round evaluation framework to measure and categorize sycophancy.
Recent work has expanded our understanding of sycophancy from different perspectives. cheng2025social identified a form called “social sycophancy”, where models avoid providing feedback that might hurt users’ feelings or self-image. Meanwhile, (zhao2024towards) demonstrated that sycophantic patterns emerge in vision-language models when processing visual content alongside user commentary.
Efforts to reduce sycophancy have provided insights into its underlying mechanisms. Synthetic data interventions can teach models to resist user pressure and maintain factual accuracy (wei2023simple), while targeted fine-tuning like pinpoint tuning addresses internal representations that drive sycophantic responses (chen2024yes). sicilia2024accounting addressed how models inappropriately mirror user confidence levels, developing uncertainty-based methods to help models express their own epistemic doubt. Other work explored steering vectors and contrastive activation methods to manipulate activations responsible for sycophancy (panickssery2023steering), revealing that sycophancy emerges from specific neural activity patterns that can be identified and modified.
@@
Mechanistic interpretability (MI) aims to reverse-engineer neural networks into human-interpretable algorithms, moving beyond traditional explainable AI focused on input-output relationships (zhang2025eap; hu2024hopfieldian; bereska2024mechanistic; kastner2024explaining; yao2025understanding; su2025understanding; zhang2025mechanistic; wang2025pahq; zhang2025understanding; dong2025understanding; yang2025understanding). Key techniques for transformers include logit-lens analysis (nostalgebraist2020interpreting), which extracts meaningful token predictions from intermediate layers, revealing what the model “believes” after each step and how these distributions converge toward the final output (stolfo2023mechanistic; zhang2025modalities). Activation patching provides a causal perspective by substituting activations between inputs to identify which components are necessary and sufficient for specific behaviors (wang2022interpretability; zhang2024locate; hong2024dissecting).
Recent applications of these MI techniques to sycophantic behavior have begun to illuminate the internal mechanisms involved, though with limitations. yu2023characterizing used head attribution to identify attention heads that resolve conflicts between memorized facts and contradictory contextual information, but their findings on factual recall (e.g., world capitals) show limited generalizability across different knowledge domains. Similarly, steering vector approaches have identified specific activation space directions corresponding to sycophantic versus truthful responding and can successfully modify model behavior through targeted interventions (panickssery2023steering), but these methods focus on controlling sycophancy rather than explaining why these particular directions emerge or what computational processes give rise to the observed activation patterns.
@@
Following previous studies, “sycophancy” in LLMs is defined as the model’s tendency to conform to a user’s explicitly stated opinion, even when that opinion is incorrect (denison2024sycophancy). To understand how models process conflicting information when user opinions contradict learned knowledge, we designed a simple experimental framework that avoids multi-stage benchmark complexity (fanous2025syceval). Prior work shows that incorrect user opinions reliably trigger sycophancy (sharma2023towards), so we use this as our primary sycophancy trigger across seven model families.
@@
Models. We evaluate seven LLMs from different families: Llama3.1 8B-Instruct (dubey2024llama), Qwen2.5 7B-Instruct (team2024qwen2; qwen2.5), OPT 6.7B (zhang2022opt), Mistral 7B-Instruct v0.3 (jiang2023mistral), Falcon 7B (tii2023falcon), OLMoE 1B-7B Instruct (muennighoff2024olmoe), and Pythia 6.9B (biderman2023pythia), all of comparable parameter size, ensuring observed differences are not due to scale effects.
Dataset. We use the Massive Multitask Language Understanding (MMLU) benchmark (hendrycks2020measuring). Its multiple-choice format provides clear ground-truth labels, while its 57 academic subjects ensure broad generalizability across knowledge domains. In our case, sycophancy occurs when the model selects the same answer choice that the user claims to believe, which we designed to always be incorrect.
@@
Evaluation Metric. For each sample, we log the model’s selected answer, and then compute three metrics: (1) sycophancy rate, or agreement rate (malmqvist2024sycophancy), the proportion of samples where the model agrees with the user’s belief; (2) accuracy, the proportion where it selects the correct answer; and (3) independent error rate, the proportion of incorrect answers that disagree with both the user’s belief and the ground truth, indicating autonomous errors.
@@
Figure 3: Responses breakdown by user expertise level. Results show expertise has negligible impact on sycophancy rates. A detailed table is at Table 2 in the Appendix.
@@
Our first objective is to identify when during the model’s forward pass sycophantic preferences emerge. Since different transformer layers encode different types of information, with later layers typically handling task-specific reasoning (tenney2019bert; rogers2021primer), we hypothesize that sycophancy arises at a specific computational stage where user opinion overrides LLM’s learned knowledge.
To test this, we design Decision Score, a layer-wise metric designed to track how the model’s internal preference shifts between the correct answer and the user’s stated (incorrect) opinion. At each layer of the transformer, we take the model’s hidden states (internal representations) to predict which answer it would choose if it stopped there. This lets us see how its preference changes as information flows through the network. Doing this gives us the prediction scores (called logits) for each of the four multiple-choice options: $l_{A},l_{B},l_{C},l_{D}$ via logit-lens (nostalgebraist2020interpreting) (more details in the Appendix). For any candidate option $x\in\{A,B,C,D\}$, we define a normalized score:
@@
Figure 4: Layered decision scores of the correct and chosen wrong answers under Plain and Opinion-only on Llama3.1 8B-Instruct. Result for Qwen can be found in Figure 11 in the Appendix.
@@
Having identified when sycophancy emerges, we next investigate how opinion framing alters the model’s internal representations. Prior work has demonstrated that different prompting strategies can lead to measurably different activation patterns (wei2022chain; hendel2023context), suggesting that semantic framing effects should be detectable in hidden state distributions.
@@
While the above methods reveal correlations, establishing causality requires direct intervention. Activation patching tests whether specific internal changes are necessary and sufficient for a behavior (meng2022locating; yeo2024towards; zhang2023towards). If our observed representational shifts truly drive sycophancy, then selectively modifying these activations should predictably alter the output.
Figure 6: Causal effects of activation patching, for both models, at respective turning-point layers found in KL divergence. (a) Plain $\rightarrow$ Opinion-only : Patching suppresses sycophancy (orange ↓) and improves accuracy (green ↑); (b) Opinion-only $\rightarrow$ Plain : Patching induces sycophancy (orange ↑) while reducing accuracy (green ↓).
@@
Figure 7: PCA projection of prompt token hidden states from Layer 32 of Llama3.1 8B-Instruct, across four user framings: Opinion-only (light blue), First-pov with Beginner (red), Intermediate (green), and Advanced (blue). Result for Qwen cna be found at Figure 12 in the Appendix.
@@
To investigate this, we turn to cognitive science, where research shows that narrative point-of-view fundamentally shapes human perception (wallace2016impact). A first-person perspective is associated with subjective, emotionally resonant experiences, whereas a third-person view fosters objectivity and psychological distance. Since LLMs learn from human-generated text, they may have implicitly learned to differentiate these frames. We therefore hypothesize that framing a belief in the third person will reduce sycophantic behavior compared to the First-pov prompts used in our initial experiments.
@@
Figure 8: 7 models consistently exhibit more sycophancy in First-pov than in Third-pov , with an average incrase of 13.6%, on all expertise levels (full result at Figure 14 in the Appendix).
As shown in Figure 8, our experimental results reveal a consistent pattern: First-pov induces higher sycophancy rates than Third-pov . A plausible explanation of this behaviour, rooted in the cognitive science principles discussed previously (wallace2016impact), is that the first-person pronoun “I” is interpreted by the model as a direct, subjective appeal from the user. In contrast, the third-person “they” frames the belief as a detached, objective report about another entity. This psychological distance appears to reduce the pressure to conform, allowing the model’s internal knowledge to influence its final output more freely.
@@
To find mechanistic evidence for the behavioral differences observed, we investigated where the model’s representations diverge under these different person framings. Specifically, using layer-wise KL divergence, we measured the difference between the hidden state distributions of First-pov and Third-pov prompts against a Plain baseline.